A coffee merchant combines coffee costing $4.50 per pound with coffee costing $3.50 per pound. How many pounds of each should be used to make 40 lb of a blend costing $3.80 per pound?

Answer by jorel1380(3719)

40x3.80=152
m+n=40
4.50m+3.50n=152
3.50m+3.50n=140
m=12
n=28
You need 12 lbs of $4.50 coffee and 28 lbs. of $3.50 coffee for your mixture..